OpenStudy (anonymous):

solve the inequality |(n)/(n+1)-1|<0.01

OpenStudy (anonymous):

Hint : \[\frac{n}{n+1}-1=\frac{n}{n+1}-\frac{n+1}{n+1}=-\frac{1}{n+1}\]

OpenStudy (anonymous):

abs cancels the negative

OpenStudy (anonymous):

i'm stuck with solving for <0.01

OpenStudy (anonymous):

better to reverse it from first step\[|\frac{1}{n+1}|<\frac{1}{100}\]\[|n+1|>100\]now solve for n

OpenStudy (anonymous):

i can see how to come up with the answer of n = 99 but i'm trying to work out the reasoning for the answer given at the value of 101

OpenStudy (anonymous):

that answer given does include 99

OpenStudy (anonymous):

conider the equation\[|x|>a\]where \(a\) is a positive real number... solution of this inequality is \(x>a\) and \(x<-a\)

OpenStudy (anonymous):

forgive me if my question is stupid, in being absolute value i am aware of the rule that you have just stated, but is there working to show

OpenStudy (anonymous):

i should also state that the answer is given as n>99 or n<-101, i think i just answered my own question

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

do u wanna know how can we prove that rule?

OpenStudy (anonymous):

yes if you don't mind

OpenStudy (anonymous):

well consider 2 cases if \(x>0\) then \(|x|=x\) and \(x>a\) if \(x<0\) then \(|x|=-x\) and \(-x>a\) multiply by \(-1\) ----> \(x<-a\)

OpenStudy (anonymous):

i'd say that's pretty much the proven

OpenStudy (anonymous):

you got a new fan

OpenStudy (anonymous):

lol...thank u

OpenStudy (anonymous):

and a similar rule : when \(|x|<a\) where \(a\) is a positive real number... then u can prove : \(-a<x<a\)

OpenStudy (anonymous):

ok hang on i gotta take all this down and go over it a few times to make it sink in, i know these rules will be handy for the future, i got a lot of catching up to do here in calculus

OpenStudy (anonymous):

i still can't find how to show the value of 101

OpenStudy (anonymous):

oh sorry i was out

OpenStudy (anonymous):

i mean -101

OpenStudy (anonymous):

we got \(|n+1|>100\) right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well according to the first rule ... \(n+1>100\) or \(n+1<-100\) from this n+1>100 ----> n>99 n+1<-100 -----> n<-101

OpenStudy (anonymous):

now i can see where the rule fits,

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks again, you'eve really taken the timer to explain things, i got another one that i have been working on its a bit harder than that one should i post it or can i just give it to you here

OpenStudy (anonymous):

np...give it here

OpenStudy (anonymous):

|(2n^2)/(n^2+4)(-2)|

OpenStudy (anonymous):

\[|\frac{2n^2}{(n^2+4)(-2)}|\]?

OpenStudy (anonymous):

the -2 is separate its not part of the denominator

OpenStudy (anonymous):

\[\left| \frac{2n^2}{(n^2+4)}(-2) \right|\]

OpenStudy (anonymous):

thats better

OpenStudy (anonymous):

u can get constants out of abs whatever it is so i can write\[\left| \frac{2n^2}{(n^2+4)}(-2) \right|=4\left| \frac{n^2}{(n^2+4)} \right|\]

OpenStudy (anonymous):

in better words : \[\left| \frac{2n^2}{(n^2+4)}(-2) \right|=|-4|\left| \frac{n^2}{(n^2+4)} \right|=4\left| \frac{n^2}{(n^2+4)} \right|\]

OpenStudy (anonymous):

because we know that \(|ab|=|a||b|\)

OpenStudy (anonymous):

i can honestly say that i have no idea where this is leading

OpenStudy (anonymous):

i'll tell u what they mean by this question... is it ok till there?

OpenStudy (anonymous):

yes but because i've never done this before i'm frozen here, i'm entirely dependant on your lead here

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

im going to prove\[\frac{n^2}{n^2+4}>0\]for all real numbers \(n\)

OpenStudy (anonymous):

can u tell me why?

OpenStudy (anonymous):

because x>a

OpenStudy (anonymous):

where a is a positive real number

OpenStudy (anonymous):

then |x|=x, yes , no|?|

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if we prove that then we can omit the abs sign

OpenStudy (anonymous):

i thought i could understand at least that much

OpenStudy (anonymous):

lets go to work on that \[\frac{n^2}{n^2+4}>0\]

OpenStudy (anonymous):

im trying but my paper is getting thin from rubbing out the mistakes

OpenStudy (anonymous):

\(n^2\) and \(n^2+4\) are positive values... so that fraction will be positive also

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

so the final result will be\[\left| \frac{2n^2}{(n^2+4)}(-2) \right|=|-4|\left| \frac{n^2}{(n^2+4)} \right|=4\left| \frac{n^2}{(n^2+4)} \right|=\frac{4n^2}{n^2+4}\]

OpenStudy (anonymous):

Mmm this is a hard one

OpenStudy (anonymous):

but u have it now ;)

OpenStudy (anonymous):

i,m still trying to find my way to the answer given at n>sqrt396 or n<-sqrt396

OpenStudy (anonymous):

because i don,t know how to deal with the quadratic

OpenStudy (anonymous):

so what is the original question?

OpenStudy (anonymous):

i think u missed something in the question

OpenStudy (anonymous):

solve the following inequalities

OpenStudy (anonymous):

is all that is missing

OpenStudy (anonymous):

u gave me \[\left| \frac{2n^2}{(n^2+4)}(-2) \right|\] well greater or less than what?

OpenStudy (anonymous):

oops <0.01

OpenStudy (anonymous):

so now the problem is\[\ \frac{4n^2}{(n^2+4)}<0.01\]

OpenStudy (anonymous):

sorry i been doing math since 8 o clock this morning

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

reverse both sides of it what will u get?

OpenStudy (anonymous):

i would only be guessing ,

OpenStudy (anonymous):

i hate to butt in, and i hope i am wrong, but i am going to bet that it was \[\left| \frac{2n^2}{(n^2+4)}-2 \right|\] and not \[\left| \frac{2n^2}{(n^2+4)}(-2) \right|\]

OpenStudy (anonymous):

like i said, i really hope i am wrong, because this is such good work

OpenStudy (anonymous):

i am thinking that you are trying to prove, using the definition, that the limit as x goes to zero is 2 that would give the inequality i wrote above

OpenStudy (anonymous):

u r right ...thats it :)

OpenStudy (anonymous):

then it is not a multiplication, but a subtraction, and the algebra will be different

OpenStudy (anonymous):

you are going to get \[|\frac{-8}{n^2+4}|<0.1\] which is much easier

OpenStudy (anonymous):

i would really like to see how you found -8

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

\[\frac{2n^2}{n^2+4}-2=\frac{2n^2-2(n^2+4)}{n^2+4}\]

OpenStudy (anonymous):

when you wrote \[\left| \frac{2n^2}{(n^2+4)}(-2) \right|\] @mukushla understandably thought it was a product because of the parentheses around the \(-2\) but i think in fact it should have been \[\left| \frac{2n^2}{(n^2+4)}-2 \right|\] which is entirely different

OpenStudy (anonymous):

i have obviously been to much math today, satellite looks like your running with it

OpenStudy (anonymous):

wait that is wrong scratch that

OpenStudy (anonymous):

you are trying to prove \[\lim_{n\to \infty}\frac{2n^2}{n^2+4}=2\]

OpenStudy (anonymous):

no my limit questions are coming up next, we are just trying to solve the inequality have you got any more input

OpenStudy (anonymous):

you are being gently led it to that topic. this is the algebra you will have to do when you get there

OpenStudy (anonymous):

great lead on

OpenStudy (anonymous):

\[|\frac{-8}{n^2+4}|<0.1\] \[|\frac{8}{n^2+4}|<0.1\] \[-0.1<\frac{8}{n^2+4}<.1\] gives two inequalities to solve, just like the first one \[-.1<\frac{8}{n^2+4}\] and \[\frac{8}{n^2+4}<.1\] normally you cannot multiply with a variable, because you do not know if it is positive or negative but in this case since \(n^2+4>0\) you can solve by multiplying both sides by \(n^2+4\)

OpenStudy (anonymous):

do you think we can solve this, because i cant keep my eyes open much longer

OpenStudy (anonymous):

ok i think solving for n gives 96, obviously i'm wrong there or missing a step

hartnn (hartnn):

because its 0.01 not 0.1........